OpenStudy (anonymous):

Need help: Find the exact sum of the following convergent series: Sum from n=0 to infinity of (2^n)/(3^n*n!)

OpenStudy (sirm3d):

\[\large \sum_{n=0}^{+\infty} \frac{ 2^n }{ 3^n n! }=\sum_{n=0}^{+\infty}\frac{ (2/3)^n }{ n! }\]what function with a Maclaurin series representation looks like the one above?